OpenStudy (anonymous):

Keisha bought a table on sale for $326.40 . This price was 66% less than the original price. What was the original price?

OpenStudy (lilg132):

100/34 = 2.941176470588235 2.941176470588235 * 326.40 = 960

OpenStudy (anonymous):

Why the precision? Isn't it more important to understand that Keisha got the table for 1/3 of the regular price? In other words the regular price is 3 times the price she paid.

OpenStudy (lilg132):

because not all questions will be that easy they need to understand the method of how to work it out

OpenStudy (anonymous):

But it is not intuitively obvious why (or event that) 100/34*price paid = original price. It seems we are encouraging the student to apply a rote formula the problem.